# A torsion theoretic interpretation for sheaves of modules and Grothendieck topologies on directed categories

### v1

@@ Abstract.
We prove that every Grothendieck topology induces a hereditary torsion pair in the category of presheaves of modules on a ringed site, and obtain a homological characterization of sheaves of modules: a presheaf of modules is a sheaf of modules if and only if it is saturated with respect to torsion presheaves, or equivalently, it is right perpendicular to torsion presheaves in the sense of Geigle and Lenzing. We also study Grothendieck topologies on directed categories ${\mathscr{C}}$, and show that every Grothendieck topology on ${\mathscr{C}}$ is a subcategory topology if and only if ${\mathscr{C}}$ is an artinian EI category. Consequently, in this case every sheaf category is equivalent to the presheaf category over a full subcategory of ${\mathscr{C}}$. Finally, we give a few potential applications of these results in group representation theory.
@@ 1. Introduction
Sheaves of modules on ringed sites have been widely applied in algebraic geometry, algebraic topology, and geometric representation theory. Recently, they were also used to investigate modular representation theory of finite groups [3, 23, 24, 25] and continuous representations of topological groups [7]. This new approach combines ideas and methods from topos theory, representation theory of categories, and representation theory of groups, and hence not only proposes a comprehensive technique to study representations of groups, but also reveals deep relations among distinct areas. A classical example is Artin’s theorem [17, III. 9, Theorems 1 and 2], from which it follows that the category of continuous representations of a topological group $G$ is equivalent to the category of sheaves of modules over a ringed site whose underlying category is an orbit category of $G$ equipped with the atomic Grothendieck topology.
@@
Let us give some details. Let ${\mathscr{C}}$ be a small category, $J$ a Grothendieck topology on the opposite category ${\mathscr{C}^{\mathrm{op}}}$, [^1] In this paper we consider representations of ${\mathscr{C}}$; that is, covariant functors from ${\mathscr{C}}$ to a concrete category. Since they are precisely presheaves over ${\mathscr{C}^{\mathrm{op}}}$, we impose Grothendieck topologies on ${\mathscr{C}^{\mathrm{op}}}$ rather than on ${\mathscr{C}}$. and $\mathcal{O}$ a structure sheaf over the site $\boldsymbol{{\mathscr{C}}}^{{\mathrm{op}}}=({\mathscr{C}^{\mathrm{op}}},J)$. Given a presheaf $V$ of modules over the ringed site $(\boldsymbol{{\mathscr{C}}}^{{\mathrm{op}}},\mathcal{O})$ (also called an $\mathcal{O}$-module) as well as an object $x$ in ${\mathscr{C}}$, we define an element $v\in V_{x}$, the value of $V$ on $x$, to be $J$-torsion if there is a covering sieve $S\in J(x)$ such that every morphism in $S$ sends $v$ to 0. The stability axiom of Grothendieck topologies guarantees that all $J$-torsion elements in $V$ form an $\mathcal{O}$-submodule; see Lemma 3.2. Consequently, the operation to take the maximal $J$-torsion submodule gives rise to a left exact endo-functor $\mathcal{T}_{J}$ on $\mathcal{O}\operatorname{-Mod}$, the category of $\mathcal{O}$-modules. We call this functor the torsion functor and say that $V$ is $J$-torsion (resp., $J$-torsion free) if $\mathcal{T}_{J}(V)=V$ (resp., $\mathcal{T}_{J}(V)=0$). If $V$ is torsion free and $\mathrm{R}^{1}\mathcal{T}_{J}(V)=0$, where $\mathrm{R}^{1}\mathcal{T}_{J}$ is the first right derived functor of $\mathcal{T}_{J}$, then we call it a $J$-saturated module.
@@
It is direct to check that $J$-saturated $\mathcal{O}$-modules are precisely $\mathcal{O}$-modules right perpendicular to $J$-torsion $\mathcal{O}$-modules in the sense of Geigle and Lenzing [11]; for details, see [11, Section 1]. Consequently, applying [11, Propositions 2.2 and 2.5], one immediately deduces the following result:
@@
Consequently, when ${\mathscr{C}}$ is a noetherian EI category, questions concerning ${\mathrm{Sh}}(\boldsymbol{{\mathscr{C}}}^{{\mathrm{op}}},\mathcal{O})$ can be completely transferred to questions about representations of full subcategories of ${\mathscr{C}}$, so we have no “generic” sheaf theory in this situation.
@@
  • skeletal full subcategories of the category ${\mathrm{FI}}$ of finite sets and injections, where ${\mathrm{FI}}$ is equivalent to the opposite category of an orbit category of the permutation group on ${\mathbb{N}}$ ([12, Examples A2.1.11(h) and D3.4.10]);
  • skeletal full subcategories of the opposite category of the category $\mathrm{FS}$ of finite sets and surjections, where $\mathrm{FS}$ is equivalent to an orbit category of the automorphism group of the free Boolean algebra on a countable infinity of generators, or the self-homeomorphism group of the Cantor set ([12, Example D3.4.12]);
@@
This paper is organized as follows. In Section 2 for the reader’s convenience we provide some background knowledge on Grothendieck topologies and sheaf theory. In Section 3 we construct a torsion theory induced by a rule $J$ satisfying partial properties of Grothendieck topology. Indeed, for most results described in this section we do not require $J$ to be a Grothendieck topology, so the reader can understand the role of each individual property. The main content of Section 4 consists of a proof of Theorem 1.1 and several consequences. Directed categories are introduced in Section 5, where we also give a proof of Theorem 1.4. In Section 6 we classify Grothendieck topologies on ${\mathscr{C}^{\mathrm{op}}}$ when ${\mathscr{C}}$ is an EI category of type ${\mathbb{N}}$. In the last section we describe some potential applications of previously established results in group representation theory, and in particular, provide a sheaf theoretic viewpoint of Alperin’s weight conjecture.
@@ 2.1. Grothendieck topologies
Throughout this paper we fix $J$ to be a rule assigning to each object $x$ a collection $J(x)$ of left ideals of morphisms on $x$. The following definition is taken from [17, III.2, Definition 1], which defines a Grothendieck topology on ${\mathscr{C}^{\mathrm{op}}}$ if $J$ satisfies all three axioms. Here we use the dual version, since we mainly consider representations of ${\mathscr{C}}$ which are covariant functors, rather than presheaves which are contravariant functors.
@@
The following fact is well known; see [17, Section III.2]. For the convenience of the reader, we give a complete proof.
@@
Note that the empty subfunctor $\emptyset$ can never belong to $J_{d}(x)$ for $x\in\operatorname{Ob}({\mathscr{C}})$. Furthermore, if ${\mathscr{C}}$ satisfies the Ore condition, then $J_{d}=J_{a}$.
@@ 2.2. Sheaf theory
Fix a Grothendieck topology $J$ on ${\mathscr{C}^{\mathrm{op}}}$ and let $\boldsymbol{{\mathscr{C}}}^{{\mathrm{op}}}=({\mathscr{C}^{\mathrm{op}}},J)$ be the corresponded site. The following definitions are taken from [17, III.4].
@@
Now we give definitions of ringed sites and sheaves of modules on ringed sites. For more details, please refer to [21, Chapter 18].
@@ 4. A torsion theoretic characterization of sheaves
Suppose that $V$ is $J$-torsion free. It has an injective hull $E$ in $\mathcal{O}\operatorname{-Mod}$ since $\mathcal{O}\operatorname{-Mod}$ is a Grothendieck category. We claim that $E$ is torsion free. Otherwise, we have $\mathcal{T}_{J}(V)\neq 0$. But $V\cap\mathcal{T}_{J}(V)=0$ since $V$ is $J$-torsion free. This is impossible because the inclusion $V\to E$ is essential by [21, Definition 47.2.1]. Thus $E$ is $J$-torsion free and injective. Applying $\operatorname{Hom}_{\mathcal{O}\operatorname{-Mod}}(T,-)$ to the short exact sequence $0\to V\to E\to E/V\to 0$ of $\mathcal{O}$-modules we deduce that $V$ is right perpendicular to every $J$-torsion module $T$ if and only if $E/V$ is $J$-torsion free. But this is equivalent to the condition that $V$ is $J$-saturated via applying the functor $\mathcal{T}_{J}$ to the short exact sequence.
$(1)\Rightarrow(2)$. Let $V$ be a sheaf of modules on $(\boldsymbol{{\mathscr{C}}}^{{\mathrm{op}}},\mathcal{O})$. It is $J$-torsion free. Since the category of $J$-torsion modules is a localizing subcategory by [11, Proposition 2.5], applying [11, Proposition 2.2] we obtain a short exact sequence
@@
We need to show the existence of amalgamations. By [17, III.5, Lemma 2], the canonical map $\eta_{V}:V\to V^{+}$ is injective where $V^{+}=V^{\sharp}$ is the sheafification of $V$. Thus we obtain the following short exact sequence in $\mathcal{O}\operatorname{-Mod}$:
@@
Furthermore, by [17, III.5, Lemma 5], $V^{+}$ is a sheaf of modules on $(\boldsymbol{{\mathscr{C}}}^{{\mathrm{op}}},\mathcal{O})$, and hence a $J$-saturated $\mathcal{O}$-module by the conclusion of the only if direction. Applying $\mathcal{T}_{J}$ to this short exact sequence, we deduce that $W$ is $J$-torsion free since $V$ is $J$-saturated.
@@
The identification follows from Theorem 4.2  [^3] We shall remind the reader that the abelian structure on $\mathcal{O}\operatorname{-Mod}^{{J\text{-}\mathrm{sat}}}$ is not inherited from the abelian structure on $\mathcal{O}\operatorname{-Mod}$. Indeed, Given a monomorphism $\alpha:V\to W$ between two $J$-saturated modules, its cokernel in $\mathcal{O}\operatorname{-Mod}$ might not coincide with the cokernel in $\mathcal{O}\operatorname{-Mod}^{{J\text{-}\mathrm{sat}}}$., while the equivalence follows from [11, Proposition 2.2(d)]. ∎
@@
which is exact. Restricting it to the full subcategory ${\mathrm{Sh}}(\boldsymbol{{\mathscr{C}}}^{{\mathrm{op}}},\mathcal{O})$, we get a functor $\Gamma_{x}$ which is only left exact since the abelian structure of ${\mathrm{Sh}}(\boldsymbol{{\mathscr{C}}}^{{\mathrm{op}}},\mathcal{O})$ is not inherited from that of $\mathcal{O}\operatorname{-Mod}$. Thus one can define $\mathrm{R}^{i}\Gamma_{x}(V)$, the $i$-th right derived functor of $\Gamma_{x}$, called the $i$-th sheaf cohomology group of $V$ at $x$. For details, please refer to [21, I.21.2].
@@
This is essentially a combination of [12, C2.3, Lemma 2.3.3] and [12, C2.3, Corollary 2.3.4], so we only give a sketch of the proof. The functor $\phi$ induces a restriction functor
@@ 5. Grothendieck topologies on artinian EI categories
The Serre quotient on the right hand side is still complicated, so we want to find a more explicit description for some special Grothendieck topologies $J$. When $J$ is a subcategory topology (or called rigid topology), this Serre quotient is equivalent to $\mathcal{O}_{\mathscr{D}}\operatorname{-Mod}$, where $\mathcal{O}_{\mathscr{D}}$ is the restriction of $\mathcal{O}$ to a full subcategory $\mathscr{D}$ of ${\mathscr{C}}$, [^4] Since $\mathcal{O}$ is a covariant functor from $\mathcal{C}$ to the category of associative rings, its shall be restricted to ${\mathscr{D}}$ rather than ${\mathscr{D}}^{{\mathrm{op}}}$. and hence becomes much more transparent. Furthermore, it has been pointed out in [2, IV.9] that all Grothendieck topologies on a finite category are rigid, and this result also holds for artinian posets (see [16, Theorem 2.12]).
@@ 5.1. Definitions and the main result
Now we introduce the rigid topology and the main theorem of this section. Let ${\mathscr{C}}$ be a small skeletal category, and $J$ a Grothendieck topology on ${\mathscr{C}^{\mathrm{op}}}$. Recall from [12, C2.2] that $x\in\operatorname{Ob}({\mathscr{C}})$ is called $J$-irreducible if $J(x)=\{{\mathscr{C}}(x,-)\}$. For $y\in\operatorname{Ob}({\mathscr{C}})$, let $S_{y}$ be the sieve generated by the family of morphisms (in ${\mathscr{C}}$ rather than ${\mathscr{C}^{\mathrm{op}}}$) from $y$ to all $J$-irreducible objects. It can be explicitly described as follows
@@
A rigid Grothendieck topology $J$ is also called a subcategory topology with respect to the full subcategory consisting of $J$-irreducible objects. When $J$ is a rigid topology, $S_{y}$ is the unique minimal covering sieve in $J(y)$, so $J(y)=\{S\subseteq{\mathscr{C}}(y,-)\mid S\supseteq S_{y}\}$. For details, see [12, C2.2, Example 2.2.4(d), Remark 2.2.5, and Definition 2.2.18].

### v2

@@ Abstract.
We prove that every Grothendieck topology induces a hereditary torsion pair in the category of presheaves of modules on a ringed site, and obtain a homological characterization of sheaves of modules: a presheaf of modules is a sheaf of modules if and only if it is saturated with respect to torsion presheaves, or equivalently, it is right perpendicular to torsion presheaves in the sense of Geigle and Lenzing. We also study Grothendieck topologies on directed categories ${\mathscr{C}}$, and show that every Grothendieck topology on ${\mathscr{C}}$ is a subcategory topology if and only if ${\mathscr{C}}$ is an artinian EI category. Consequently, in this case every sheaf category is equivalent to the presheaf category over a full subcategory of ${\mathscr{C}}$. Finally, we classify all Grothendieck topologies on a special type of noetherian EI categories, and extend the locally self-injective property of representations of ${\mathrm{FI}}$ and ${\mathrm{VI}}$ to representations of their infinite full subcategories. Some potential applications in group representation theory are given at the end of this paper.
@@ 1. Introduction
Sheaves of modules on ringed sites have been widely applied in algebraic geometry, algebraic topology, and geometric representation theory. Recently, they were also used to investigate modular representation theory of finite groups [3, 25, 26, 27] and continuous representations of topological groups [7]. This new approach combines ideas and methods from topos theory, representation theory of categories, and representation theory of groups, and hence not only proposes a comprehensive technique to study representations of groups, but also reveals deep relations among distinct areas. A classical example is Artin’s theorem [18, III. 9, Theorems 1 and 2], from which it follows that the category of continuous representations of a topological group $G$ is equivalent to the category of sheaves of modules over a ringed site whose underlying category is an orbit category of $G$ equipped with the atomic Grothendieck topology.
@@
Let us give some details. Let ${\mathscr{C}}$ be a small category, $J$ a Grothendieck topology on the opposite category ${\mathscr{C}^{\mathrm{op}}}$, [^1] In this paper we consider representations of ${\mathscr{C}}$; that is, covariant functors from ${\mathscr{C}}$ to a concrete category. Since they are precisely presheaves over ${\mathscr{C}^{\mathrm{op}}}$, we impose Grothendieck topologies on ${\mathscr{C}^{\mathrm{op}}}$ rather than on ${\mathscr{C}}$. and $\mathcal{O}$ a structure sheaf over the site $\boldsymbol{{\mathscr{C}}}^{{\mathrm{op}}}=({\mathscr{C}^{\mathrm{op}}},J)$. Given a presheaf $V$ of modules over the ringed site $(\boldsymbol{{\mathscr{C}}}^{{\mathrm{op}}},\mathcal{O})$ (also called an $\mathcal{O}$-module) as well as an object $x$ in ${\mathscr{C}}$, we define an element $v\in V_{x}$, the value of $V$ on $x$, to be $J$-torsion if there is a covering sieve $S\in J(x)$ such that every morphism in $S$ sends $v$ to 0. The axioms of Grothendieck topologies guarantee that all $J$-torsion elements in $V$ form an $\mathcal{O}$-submodule; see Lemma 3.2. Consequently, the operation to take the maximal $J$-torsion submodule gives rise to a left exact endo-functor $\mathcal{T}_{J}$ on $\mathcal{O}\operatorname{-Mod}$, the category of $\mathcal{O}$-modules. We call this functor the torsion functor and say that $V$ is $J$-torsion (resp., $J$-torsion free) if $\mathcal{T}_{J}(V)=V$ (resp., $\mathcal{T}_{J}(V)=0$). If $V$ is torsion free and $\mathrm{R}^{1}\mathcal{T}_{J}(V)=0$, where $\mathrm{R}^{1}\mathcal{T}_{J}$ is the first right derived functor of $\mathcal{T}_{J}$, then we call it a $J$-saturated module.
@@
It is direct to check that $J$-saturated $\mathcal{O}$-modules are precisely $\mathcal{O}$-modules right perpendicular to $J$-torsion $\mathcal{O}$-modules in the sense of Geigle and Lenzing [12]; for details, see [12, Section 1]. Consequently, applying [12, Propositions 2.2 and 2.5], one immediately deduces the following result:
@@
Consequently, when ${\mathscr{C}}$ is a noetherian EI category, questions concerning ${\mathrm{Sh}}(\boldsymbol{{\mathscr{C}}}^{{\mathrm{op}}},\mathcal{O})$ can be completely transferred to questions about representations of full subcategories of ${\mathscr{C}}$, so we have no “generic” sheaf theory in this situation. Explicitly, by the Comparison Lemma [13, C2.2, Theorem 2.2.3] (see also [13, C2.2, Example 2.2.4(d)]), one has:
@@
  • skeletal full subcategories of the category ${\mathrm{FI}}$ of finite sets and injections, where ${\mathrm{FI}}$ is equivalent to the opposite category of an orbit category of the permutation group on ${\mathbb{N}}$ ([13, Examples A2.1.11(h) and D3.4.10]);
  • skeletal full subcategories of the opposite category of the category $\mathrm{FS}$ of finite sets and surjections, where $\mathrm{FS}$ is equivalent to an orbit category of the automorphism group of the free Boolean algebra on a countable infinity of generators, or the self-homeomorphism group of the Cantor set ([13, Example D3.4.12]);
@@
We can use Theorem 1.6 and the Comparison Lemma to deduce some interesting results in representation theory of combinatorial categories. For instance, let $k$ be a field of characteristic 0, and let ${\mathscr{C}}$ be the category of finite sets and injections, or the category of finite dimensional vector spaces over a finite field $\mathbb{F}_{q}$ and $\mathbb{F}_{q}$-linear injections. It has been proved in [11] (see also [22]) that every indecomposable projective representation of ${\mathscr{C}}$ over $k$ is also injective. It turns out this locally self-injective property holds for all (essentially) infinite full subcategories of ${\mathscr{C}}$.
Theorem 1.7.
Let ${\mathscr{C}}$ be specified as above, $k$ a field of characteristic 0, and ${\mathscr{D}}$ a full subcategory of ${\mathscr{C}}$ containing infinitely many objects (up to isomorphism). Then every indecomposable projective ${\mathscr{D}}$-module over $k$ is also injective.
For a generalized version of this result, see Theorem 7.1.
This paper is organized as follows. In Section 2 for the reader’s convenience we provide some background knowledge on Grothendieck topologies and sheaf theory. In Section 3 we construct a torsion theory induced by a rule $J$ satisfying partial properties of Grothendieck topology. Indeed, for most results described in this section we do not require $J$ to be a Grothendieck topology, so the reader can understand the role of each individual property. The main content of Section 4 consists of a proof of Theorem 1.1 and several consequences. Directed categories are introduced in Section 5, where we also give a proof of Theorem 1.4. In Section 6 we classify Grothendieck topologies on ${\mathscr{C}^{\mathrm{op}}}$ when ${\mathscr{C}}$ is an EI category of type ${\mathbb{N}}$, proving Theorem 1.6. In the last section we describe some potential applications of previously established results and prove Theorem 1.7.
@@ 2.1. Grothendieck topologies
Throughout this paper we fix $J$ to be a rule assigning to each object $x$ a collection $J(x)$ of left ideals of morphisms on $x$. The following definition is taken from [18, III.2, Definition 1], which defines a Grothendieck topology on ${\mathscr{C}^{\mathrm{op}}}$ if $J$ satisfies all three axioms. Here we use the dual version, since we mainly consider representations of ${\mathscr{C}}$ which are covariant functors, rather than presheaves which are contravariant functors.
@@
The following fact is well known; see [18, Section III.2]. For the convenience of the reader, we give a complete proof.
@@
Note that the empty subfunctor $\emptyset$ can never belong to $J_{d}(x)$ for $x\in\operatorname{Ob}({\mathscr{C}})$. Furthermore, if ${\mathscr{C}}$ satisfies the right Ore condition, then $J_{d}=J_{a}$.
@@ 2.2. Sheaf theory
Fix a Grothendieck topology $J$ on ${\mathscr{C}^{\mathrm{op}}}$ and let $\boldsymbol{{\mathscr{C}}}^{{\mathrm{op}}}=({\mathscr{C}^{\mathrm{op}}},J)$ be the corresponded site. The following definitions are taken from [18, III.4].
@@
Now we give definitions of ringed sites and sheaves of modules on ringed sites. For more details, please refer to [23, Chapter 18].
@@ 4. A torsion theoretic characterization of sheaves
Suppose that $V$ is $J$-torsion free. It has an injective hull $E$ in $\mathcal{O}\operatorname{-Mod}$ since $\mathcal{O}\operatorname{-Mod}$ is a Grothendieck category. We claim that $E$ is torsion free. Otherwise, we have $\mathcal{T}_{J}(V)\neq 0$. But $V\cap\mathcal{T}_{J}(V)=0$ since $V$ is $J$-torsion free. This is impossible because the inclusion $V\to E$ is essential by [23, Definition 47.2.1]. Thus $E$ is $J$-torsion free and injective. Applying $\operatorname{Hom}_{\mathcal{O}\operatorname{-Mod}}(T,-)$ to the short exact sequence $0\to V\to E\to E/V\to 0$ of $\mathcal{O}$-modules we deduce that $V$ is right perpendicular to every $J$-torsion module $T$ if and only if $E/V$ is $J$-torsion free. But this is equivalent to the condition that $V$ is $J$-saturated via applying the functor $\mathcal{T}_{J}$ to the short exact sequence.
$(1)\Rightarrow(2)$. Let $V$ be a sheaf of modules on $(\boldsymbol{{\mathscr{C}}}^{{\mathrm{op}}},\mathcal{O})$. It is $J$-torsion free. Since the category of $J$-torsion modules is a localizing subcategory by [12, Proposition 2.5], applying [12, Proposition 2.2] we obtain a short exact sequence
@@
We need to show the existence of amalgamations. By [18, III.5, Lemma 2], the canonical map $\eta_{V}:V\to V^{+}$ is injective where $V^{+}=V^{\sharp}$ is the sheafification of $V$. Thus we obtain the following short exact sequence in $\mathcal{O}\operatorname{-Mod}$:
@@
Furthermore, by [18, III.5, Lemma 5], $V^{+}$ is a sheaf of modules on $(\boldsymbol{{\mathscr{C}}}^{{\mathrm{op}}},\mathcal{O})$, and hence a $J$-saturated $\mathcal{O}$-module by the conclusion of the only if direction. Applying $\mathcal{T}_{J}$ to this short exact sequence, we deduce that $W$ is $J$-torsion free since $V$ is $J$-saturated.
@@
The identification follows from Theorem 4.2  [^3] We shall remind the reader that the abelian structure on $\mathcal{O}\operatorname{-Mod}^{{J\text{-}\mathrm{sat}}}$ is not inherited from the abelian structure on $\mathcal{O}\operatorname{-Mod}$. Indeed, Given a monomorphism $\alpha:V\to W$ between two $J$-saturated modules, its cokernel in $\mathcal{O}\operatorname{-Mod}$ might not coincide with the cokernel in $\mathcal{O}\operatorname{-Mod}^{{J\text{-}\mathrm{sat}}}$., while the equivalence follows from [12, Proposition 2.2(d)]. ∎
@@
which is exact. Restricting it to the full subcategory ${\mathrm{Sh}}(\boldsymbol{{\mathscr{C}}}^{{\mathrm{op}}},\mathcal{O})$, we get a functor $\Gamma_{x}$ which is only left exact since the abelian structure of ${\mathrm{Sh}}(\boldsymbol{{\mathscr{C}}}^{{\mathrm{op}}},\mathcal{O})$ is not inherited from that of $\mathcal{O}\operatorname{-Mod}$. Thus one can define $\mathrm{R}^{i}\Gamma_{x}(V)$, the $i$-th right derived functor of $\Gamma_{x}$, called the $i$-th sheaf cohomology group of $V$ at $x$. For details, please refer to [23, I.21.2].
@@
This is essentially a combination of [13, C2.3, Lemma 2.3.3] and [13, C2.3, Corollary 2.3.4], so we only give a sketch of the proof. The functor $\phi$ induces a restriction functor
@@ 5. Grothendieck topologies on artinian EI categories
The Serre quotient on the right hand side is still complicated, so we want to find a more explicit description for some special Grothendieck topologies $J$. When $J$ is a subcategory topology (or called rigid topology), this Serre quotient is equivalent to $\mathcal{O}_{\mathscr{D}}\operatorname{-Mod}$, where $\mathcal{O}_{\mathscr{D}}$ is the restriction of $\mathcal{O}$ to a full subcategory $\mathscr{D}$ of ${\mathscr{C}}$, [^4] Since $\mathcal{O}$ is a covariant functor from $\mathcal{C}$ to the category of associative rings, its shall be restricted to ${\mathscr{D}}$ rather than ${\mathscr{D}}^{{\mathrm{op}}}$. and hence becomes much more transparent. Furthermore, it has been pointed out in [2, IV.9] that all Grothendieck topologies on a finite category are rigid, and this result also holds for artinian posets (see [17, Theorem 2.12]).
@@ 5.1. Definitions and the main result
Now we introduce the rigid topology and the main theorem of this section. Let ${\mathscr{C}}$ be a small skeletal category, and $J$ a Grothendieck topology on ${\mathscr{C}^{\mathrm{op}}}$. Recall from [13, C2.2] that $x\in\operatorname{Ob}({\mathscr{C}})$ is called $J$-irreducible if $J(x)=\{{\mathscr{C}}(x,-)\}$. For $y\in\operatorname{Ob}({\mathscr{C}})$, let $S_{y}$ be the sieve generated by the family of morphisms (in ${\mathscr{C}}$ rather than ${\mathscr{C}^{\mathrm{op}}}$) from $y$ to all $J$-irreducible objects. It can be explicitly described as follows
@@
A rigid Grothendieck topology $J$ is also called a subcategory topology with respect to the full subcategory consisting of $J$-irreducible objects. When $J$ is a rigid topology, $S_{y}$ is the unique minimal covering sieve in $J(y)$, so $J(y)=\{S\subseteq{\mathscr{C}}(y,-)\mid S\supseteq S_{y}\}$. For details, see [13, C2.2, Example 2.2.4(d), Remark 2.2.5, and Definition 2.2.18].